OpenStudy (anonymous):

Will Fan and MEdal!!!! Express the function y = 2x^2 + 8x + 1 in vertex form.

OpenStudy (academicgurusinc):

f(x) = 2x^2 + 8x + 1 Step one: factor out 2 f(x) = 2(x^2 + 4x) + 1 Step two: find perfect square for x^2 + 4x f(x) = 2(x^2 + 4x + 4) + 1 - (4*2) = 2(x^2 + 2)^ 2 -7 Therfore, the vertex form is 2(x +2)'^2 -7 vertex (-2, -7) We hope this helps. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)